. A 47 kg child is standing on a scale in an elevator, use this for each of the following cases:

A) What does the scale read while the elevator is traveling down while slowing at 2.5 m/s2?

B) What is the acceleration is the scale reads 555 N?
1
Expert's answer
2020-11-17T13:02:41-0500

(A) Let’s apply the Newton’s Second Law of Motion:


"\sum F_y=ma_y,""N-mg=-ma,"

here, "N" is the normal force directed upward, "mg" is the force of gravity directed downward, "a" is the acceleration of an elevator (since the acceleration of an elevator directed downward it will be with sign minus).

Then, we get:


"N=m(g-a)=47\ kg\cdot(9.8\ \dfrac{m}{s^2}-2.5\ \dfrac{m}{s^2})=343.1\ N."

(B) Let’s apply the Newton’s Second Law of Motion:


"\sum F_y=ma_y,""N-mg=ma,""a=\dfrac{N-mg}{m},""a=\dfrac{555\ N-47\ kg\cdot 9.8\ \dfrac{m}{s^2}}{47\ kg}=2.0\ \dfrac{m}{s^2}."

The sign plus means that the elevator moves upward.

Answer:

(A) "N=343.1\ N."

(B) "a=2.0\ \dfrac{m}{s^2}," upward.